Solve the following system of equations

dx/dt-dy/dt-2x-4y=t2

dx/dt+dy/dt-x-y=1


1
Expert's answer
2021-08-10T13:02:49-0400
"(x+y)'-(x+y)=1"

"\mu(t)=e^{-t}"

"e^{-t}(x+y)'-e^{-t}(x+y)=e^{-t}"

"(e^{-t}(x+y))'=e^{-t}"

Integrate


"\int d(e^{-t}(x+y))=\int e^{-t}dt"

"e^{-t}(x+y)=-e^{-t}+C_1"

"x+y=-1+C_1e^{t}"

"y=-x-1+C_1e^{t}"


"y'=-x'+C_1e^{t}"

"x'+x'-C_1e^{t}-2x+4x+4-4C_1e^{t}=t^2"

"2x'+2x=5C_1e^{t}+t^2-4"

"x'+x=\dfrac{5}{2}C_1e^{t}+\dfrac{1}{2}t^2-2"

"\mu(t)=e^t"

"e^{t}x'+e^{t}x=\dfrac{5}{2}C_1e^{2t}+\dfrac{1}{2}e^{t}t^2-2e^{t}"

"d(e^{t}x)=\dfrac{5}{2}C_1e^{2t}+\dfrac{1}{2}e^{t}t^2-2e^{t}"

Integrate


"\int d(e^{t}x)=\int (\dfrac{5}{2}C_1e^{2t}+\dfrac{1}{2}e^{t}t^2-2e^{t})dt"

"\int e^{t}t^2dt=t^2e^t-2te^t+2e^{t}+C_3"

"e^{t}x=\dfrac{5}{4}C_1e^{2t}+\dfrac{1}{2}e^{t}t^2-e^{t}t+e^{t}-2e^{t}+C_2"

"x(t)=\dfrac{5}{4}C_1e^{t}+\dfrac{1}{2}t^2-t-1+C_2e^{-t}"

"y(t)=-\dfrac{5}{4}C_1e^{t}-\dfrac{1}{2}t^2+t+1+C_2e^{-t}-1+C_1e^{t}"

"y(t)=-\dfrac{1}{4}C_1e^{t}-\dfrac{1}{2}t^2+t+C_2e^{-t}"